### v1

# Critical Filaments and Superconductivity in Quasiperiodic Twisted Bilayer Graphene

Abstract
Moiré materials can exhibit electronic topological features yet are inherently quasiperiodic. Nonetheless, the localizing tendency of quasiperiodicity can be prevented by topology. We consider a quasiperiodic variant of the chiral Bistritzer-MacDonald model for twisted bilayer graphene with two incommensurate moiré potentials. We observe “filaments” linking magic angles with enhanced density of states and fractal wave functions that evade localization; states away from the filaments mimic fractal surface states of dirty topological superconductors. We demonstrate that topological quasiperiodicity can broadly enhance superconductivity without magic-angle fine-tuning.
Amplifying electronic interactions by band-velocity engineering drives much of the research in moiré materials, exemplified by twisted bilayer graphene (TBLG) [1]. Adjusting the twist angle between layers tunes the carrier velocity in the moiré bands [2, 3], leading to (nearly) flat bands at certain magic angles [4, 5]. Band flattening strongly enhances interaction effects, and led to the observation of many correlated phases including superconductivity [1]. An additional feature of TBLG is the topological character of the effective bands [7, 8, 6, 9], which is enabled by the coupling of same-chirality Dirac electrons from the two layers. The confluence of strong interaction effects and topology at the magic angles facilitates the rich panorama of phenomena observed in TBLG, but also makes the experiments and the theoretical understanding thereof sensitive to fine details.
The quasiperiodic character of the microscopic structure leads to incommensurate tunneling with respect to the underlying lattices. At normal twist angles in TBLG showing correlated behavior, these effects are negligible. In contrast, incommensuration becomes more pronounced in trilayer systems where two moiré patterns can compete; superconductivity was very recently observed in such a quasiperiodic crystal [10]. Quasiperiodicity can even occur with respect to the often benign, unaligned subtrate: if twisted bilayer graphene is brought into alignment with the boron nitride substrate, one can see the emergence of quasicrystalline structure in scanning tunneling microscopy experiments [11].
In topological materials, topology competes with the localizing nature of disorder in low-dimensions, protecting boundary and bulk states in nontrivial bands. Quasiperiodicity can also drive localization [12, 13]. The interplay between weak localization and interactions can sometimes produce surprising effects. For example, the average Cooper pairing amplitude can be enhanced near a superconductor-insulator transition if long-range Coulomb interactions are externally screened. The mechanism involves the Russian-doll-like nesting (“Chalker scaling”) of multifractal single-particle wave functions near the localization transition [14, 15], which boosts interaction matrix elements. A similar enhancement in self-consistent BCS mean-field theory was observed in quasiperiodic models [16, 17, 18].
This raises an important question: can the incommensurate effects in moiré materials change the single-particle electronic wave functions in such a way that correlated behavior, such as superconductivity, can be enhanced? Here we numerically study a quasiperiodic variant of the chiral Bistritizer-MacDonald (BM) model for TBLG. One of the central results of this Letter is that incommensuration effects can robustly amplify superconducting pairing without the need to fine-tune to magic angles.
Previous work on incommensuration in Dirac and moiré materials focused on tight-binding models where the lattice provided incommensuration [19, 20, 21]; these works found single-particle, critical phases associated with “magic-angles” providing the first hint of multifractality. In contrast, the model considered here captures long wavelength physics where the moiré pattern (and the resulting super-moiré pattern) governs the low energy processes and topology emerges. We therefore expect our observations here are relevant for twisted stackings past the bilayer setup.
The original BM model for twisted bilayer graphene [2, 4] possesses an emergent particle-hole symmetry [6], and it falls into class C in the $10$-fold classification of topological matter [22] and 2D Dirac Hamiltonians [23]. If AA tunneling vanishes, an additional chiral symmetry emerges [5]. This chiral BM model takes the same form as the 2D surface theory of a bulk class CI topological superconductor (TSC) [24]; the version with only particle-hole symmetry corresponds to the same topological surface, subject to orbital time-reversal symmetry breaking.
It has long been known that quenched disorder which respects the chiral symmetry leads to remarkable results for class-CI topological Dirac surface states, including a power-law vanishing density of states (DOS) $\nu(E)\sim\left|E\right|^{1/7}$ [25] and critically Anderson delocalized low-energy wave functions with universal multifractal statistics [26]. Recent results have demonstrated that a different kind of universal wave-function criticality can extend over a wide range of the energy spectrum for these special dirty Dirac models (“spectrum-wide quantum criticality,” SWQC [27, 28]), although this can be disrupted at the 2D boundary of bulk topological systems by lattice-scale effects [29].
In this Letter, we explore the consequences of quasiperiodicity, revealing DOS enhancements and robust wave-function fractality, and demonstrate its connection to the disordered case. Using self-consistent BCS calculations [18], we show that the quasiperiodicity-induced fractality enhances superconducting pairing [14, 15] without fine-tuning to magic angles.
[…]
Model.—We consider a quasiperiodic version of the chiral Bistritzer-MacDonald (qBM) model [4, 5],
$h(\mathbf{r})=-iv_{F}\,\bm{\sigma}\cdot\nabla+V_{\mathbf{q}}(\mathbf{r})+V_{\mathbf{q}^{\prime}}(\mathbf{r})\,.$ (1)
The Pauli matrices $\sigma_{1,2}$ act on the sublattice space, and $V_{\mathbf{q}}(\mathbf{r})$ and $V_{\mathbf{q}^{\prime}}(\mathbf{r})$ denote moiré potentials coupling the two layers in AB- and BA-stacking regimes,
[…]
Here $\tau^{\pm}=(\tau_{x}\pm i\tau_{y})/2$ (and similarly for $\sigma^{\pm}$) with $\tau_{j}$ as the Pauli matrices that link the layers. The potentials $U_{0}$ and $U_{1}$ take the form [4]
$U_{0}(\mathbf{r})$ $=e^{-i\mathbf{q}_{1}\cdot\mathbf{r}}+e^{-i\mathbf{q}_{2}\cdot\mathbf{r}}+e^{-i\mathbf{q}_{3}\cdot\mathbf{r}}\,$ (3)
$U_{1}(\mathbf{r})$ $=e^{-i\mathbf{q}_{1}\cdot\mathbf{r}}+e^{-i\frac{2\pi}{3}}e^{-i\mathbf{q}_{2}\cdot\mathbf{r}}+e^{i\frac{2\pi}{3}}e^{-i\mathbf{q}_{3}\cdot\mathbf{r}}\,.$ (4)
Here $\mathbf{q}_{1}=k_{\theta}(0,1)$ and $\mathbf{q}_{2/3}=k_{\theta}(\mp\sqrt{3}/2,-1/2)$ are the reciprocal lattice vectors of the usual moiré potential. The momentum $k_{\theta}=2K\sin\theta/2$ defines the moiré scale (with $K$ as the distance from the K point to the $\Gamma$ point in the original graphene lattice). The second moiré potential $V_{\mathbf{q}^{\prime}}(\mathbf{r})$ takes the same form but with different reciprocal vectors $\{\mathbf{q}^{\prime}_{i}\}$ and potential strength $A_{2}$. When the two sets $\{\mathbf{q}_{i}\}$ and $\{\mathbf{q}^{\prime}_{i}\}$ are mutually incommensurate, we have a quasiperiodic model. In this Letter, we focus on the case $\mathbf{q}_{i}=\beta\,\mathbf{q}_{i}^{\prime}$ with $\beta$ the inverse golden ratio $(\sqrt{5}-1)/2$; we have verified that similar results obtain for other (e.g., silver) ratios [30].
The qBM model is realized in momentum space using a momentum cutoff $\Lambda=N_{\Lambda}k_{\theta}/2$. Focusing on $V_{\mathbf{q}}(\mathbf{r})$, the moiré Brillouin zone (mBZ) has linear size $k_{\theta}$, and we keep $2N_{\Lambda}^{2}$ moiré bands. Without a second potential, a momentum space lattice of linear size $N_{\Lambda}$ is sufficient; however, $\mathbf{q}^{\prime}_{i}$ allows hoppings within the mBZ. Since $|\mathbf{q}_{i}^{\prime}|=k_{\theta}/\beta$, we controllably reach larger system sizes with rational approximations of $\beta$, i.e. $\beta_{n}=F_{n}/F_{n+1}$ for Fibonacci numbers $F_{n}$. In this case, the denominator $\frac{|\mathbf{q}_{i}^{\prime}|}{k_{\theta}}=\frac{F_{n+1}}{F_{n}}$ determines the linear size of momentum space points required within one mBZ. Therefore, the full momentum space lattice has linear size $N_{\Lambda}F_{n}$ and spacing $\Delta k=k_{\theta}/F_{n}$; in the incommensurate limit ($n\rightarrow\infty$), all momentum can be accessed. In the following, we work in units such that $v_{F}=1$ and $k_{\theta}=\left|\mathbf{q}_{i}\right|=1$. Our setup is illustrated in Fig. 1.
Figure 2: The zero-energy density of states $\nu(0)$ for the qBM model in chiral limit. Here $A_{1}$ and $A_{2}$ are the strengths of the two sets of moiré potentials. (a): The map of $\nu(0)$ for the two moiré potentials. The zero-energy DOS has large peaks at the magic angles of the $A_{1}$ and $A_{2}$ potentials [indicated by red stars in (a)]. The magic angles of $A_{1}$ and $A_{2}$ are connected by several filaments, where $\nu(0)$ peaks. (b), (c) and (d): $\nu(0)$ along the cuts at $A_{1}=0.4$, $1$ and $1.5$, respectively. The peaks correspond to the filaments and $\nu(0)$ remains nonzero between these. This is our first hint that quasi-periodicity has driven a single-particle phase transition away from the semimetal in these regions. The data are obtained with KPM for a momentum lattice with $N_{\Lambda}=17$ and $M=89$. We use Chebyshev expansion order $N_{\scriptscriptstyle\mathcal{C}}=2^{11}$ for the DOS map (a) and $N_{\scriptscriptstyle\mathcal{C}}=2^{11}$, $2^{12}$ and $2^{13}$ in (b-d).
Density of states and fractal wave functions.—Dirac cones have a vanishing DOS $\nu(E)\sim|E|/v^{2}$ with velocity $v$; at the magic angle $v\rightarrow 0$ and $\nu(0)$ diverges. In the quasiperiodic case, at small $A_{1,2}$ the semimetal is stable and eigenstates remain plane-wave-like [36]. However, when $\nu(0)>0$ (or vanishes with a small power-law), multifractality–typically associated to an Anderson metal-insulator transition [37]–can occur. To reveal these (potential) single-particle phases, Fig. 2 shows $\nu(0)$ for the qBM model evaluated with the kernel polynomial method (KPM) [38]. Fig. 2(a) shows the $\nu(0)$ landscape while $A_{1,2}$ vary. There are large lakes with $\nu(0)>0$ separated by filaments where the DOS is strongly enhanced. The filaments connect the magic angles (red stars) of the monochromatic potentials; both filaments and lakes host multifractal wave functions (see below). Fig. 2(b-d) show particular cuts. In the regime close to the $A_{1}$ or $A_{2}$ axes away from the magic angles, the system remains semimetallic and $\nu(0)\rightarrow 0$. By contrast, $\nu(0)$ peaks at the filaments and either remains nonzero or vanishes with a nontrivial power-law between the filaments [30]. The system exhibits semi-metallic behavior for small $A_{1}$ or $A_{2}$ prior to the monochromatic magic angles. In the regions of comparable $A_{1}$ and $A_{2}$, between the filaments, $\nu(0)$ is generally nonzero; in fact, as we show below, the system exhibits SWQC behavior in these regions associated to class-CI surface states [27].
Figure 3: The second multifractal dimension $\tau_{2}$ averaged for states near zero energy and zero energy DOS $\nu(E=0)$ for the qBM model. (a) and (b): along the vertical cuts at $A_{1}=0.4$ and $1$, respectively; (c) and (d): along the horizontal cuts at $A_{2}=0.2$ and $A_{2}=0.5$, respectively. The navy blue dashed and orange dot-dashed lines indicate $\tau_{2}=3/2$ for the WZW states and $\tau_{2}=7/4$ for the SQHT states, respectively.
As illustrated in Fig. 1, both cutoff $\Lambda$ and number of moiré unit cells $F_{n}$ determine the momentum space lattice, which need to both be large; therefore, to determine wave functions, we employ the Chebyshev filtering and shaking method [39, 40]. This method obtains eigenfunctions in a small energy window, which could contain thousands of states near a narrow band [30]. We use the method to obtain eigenfunctions for the qBM model with $\sim 10^{6}$ lattice sites.
Figure 4: The second multifractal dimension $\tau_{2}$ versus energy and the wave functions at the one of filaments and in the lakes between the filaments. (a): $\tau_{2}$-$E$ at the filament with $A_{1}=0.4$ and $A_{2}=0.36$; (b): same as (a) but in the lakes between filaments with $A_{1}=0.4$ and $A_{2}=0.6$. (c)-(e): the wave functions at the filament for states at $E=0$, $0.2$ and $0.425$; (f)-(h): the wave functions in the lake for states at $E=0$, $0.2$ and $0.5$. The shadows in (a) and (b) show the energy gaps in the spectrum. The navy dashed and orange dot-dashed line in (a) and (b) show the reference value of $\tau_{2}$ for the WZW and SQHT states, respectively. Here we $N_{\Lambda}=11$, $M=55$ and the wave functions reside on a $605\times 605$ lattice in real space.
The statistics of single-particle wave function can be characterized by the inverse participation ratio (IPR),
$P_{q}=\sum_{i}\left|\psi_{i}\right|^{2q}\,,$ (5)
Here $\psi_{i}$ is the (normalized) wave function in real space, obtained by Fourier transform of the momentum-space wave function. For extended wave functions, $P_{q}\sim L^{-d(q-1)}$, with $L$ the linear dimension of the system; for localized wave functions, $P_{q}\sim L^{0}$. Further, if $P_{q}\sim L^{-\tau_{q}}$ and $0<\tau_{q}<d(q-1)$, that wave function is critical. In the thermodynamic limit ($L\to\infty$), the multifractal dimension can be evaluated via
$\tau_{q}=-\log{P_{q}}/\log{L}\,.$ (6)
The $\tau_{q}$ can also be obtained with the box averaged IPR, $P_{q}^{\scriptscriptstyle\mathsf{box}}=\sum_{i_{b}}\left(\sum_{\text{box}}\left|\psi_{i}\right|^{2}\right)^{q}\propto b^{\tau_{q}}\,.$ with $b$ the linear size of the box. We use both methods to calculate the $\tau_{q}$ and they are consistent with each other for the appropriate choice of box sizes.
The second multifractal dimension $\tau_{2}$ is shown in Fig. 3 for the qBM model along the cuts $A_{1}=0.4$ and $1$. Corresponding $\tau_{3}$ plots appear in [30]. Note that a value of $\tau_{2}\approx 2$ indicates ballistic wavefunctions, and Fig. 3(a) shows states like that for small $A_{2}$ and $A_{1}=0.4$. We observe a correlation between $\nu(0)$ and multifractality; at the first filament (with increasing $A_{2}$), $\tau_{2}$ decreases, fluctuates for stronger $A_{2}$, and dips near the additional filaments. The value of $\tau_{2}$ near the filaments in Fig. 3(a,c) indicates that the wave functions there are strongly multifractal, see also Fig. 4(c,d,g). Note that Fig. 3(c) follows the pattern for magic-angle semimetals as outlined in [20, 21]. In other regimes, such as between the filaments, the wave functions are also multifractal with a different value of $\tau_{2}$. In the nonperturbative regime, e.g., Fig. 3(b) along $A_{1}=1$, $\tau_{2}$ fluctuates around $3/2$ and is only partially correlated with $\nu(0)$. However, at this point $\tau_{2}$ appears to have saturated to the known value $\tau_{2}=3/2$ for critical states in class-CI Dirac models, where the emergent Wess-Zumino-Witten (WZW) term enables the zero-energy state to escape Anderson localization [26, 37, 23]. [Fig. 3(d) also illustrates this saturation.]
Finite-energy surface-state wave functions of a dirty TSC can also evade localization via SWQC, with critical statistics tied to a quantum-Hall plateau transition [27, 41, 28, 29]. For the class-CI qBM model, we anticipate finite-energy states with SWQC associated to the spin quantum-Hall transition (SQHT) [27, 28]. To compare with finite energies in the qBM model, we take two separate values, on the filament ($A_{1}=0.4$ and $A_{2}=0.36$) and between the filaments ($A_{1}=0.4$ and $A_{2}=0.6$) [see Fig 4(a,b)], and examine a range of energies, extracting $\tau_{2}$. On the filament [Fig 4(a)], the low-energy states are strongly multifractal (with $\tau_{2}\approx 1$) with a few isolated peaks [energies $E=0,0.2$ in both Figs. 4(c,d)]; the higher energies oscillate between SQHT states [$\tau_{2}\approx 7/4$ and less strongly peaked, see Fig. 4(e)] and strongly multifractal states. The filaments, originating from the flat bands at the magic-angles, have hybridized those states strongly into these multifractal states. In between the filaments [Fig 3(b)], the low-energy wave functions are less multifractal than on the filament with WZW-like states [$\tau_{2}\approx 3/2$ and $\tau_{3}\approx 5/2$, Fig. 4(g)] and the high-energy states that reside deep within moiré bands are SQHT states [$\tau_{2}\approx 7/4$ and $\tau_{3}\approx 13/4$, Fig. 4(h)]. Near the energy gaps, the wave functions return to strongly multifractal or WZW-like states.
Figure 5: The superconductivity order parameter $\Delta$, zero energy DOS $\nu(0)$, distribution of the local order parameter $\Delta_{i}$ and superfluid stiffness for the qBM-BCS model. (a): $\Delta$ and $\nu(E=0)$ along the cut $A_{1}=0.4$; (b): same as (a) along the cut $A_{1}=1$; (c): the distribution of local order parameter $\Delta_{i}$ along $A_{1}=0.4$; (d) same as (c) along the cut $A_{1}=1$; (e): superfluid stiffness (red) along $A_{1}=0.4$ (clean limit value $2\Delta/\pi$ indicated by black dashed curve); (f): same as (e) along $A_{1}=1$.
Superconductivity.—The Bogoliubov-de Gennes Hamiltonian for the qBM model with pseudospin, layer, and spin-singlet $s$-wave pairing is given by
[…]
Here $\chi$ is the Nambu spinor spanning sublattice ($\sigma$), layer ($\tau$), and particle-hole spaces ($\mu$) and
$h_{\mathsf{BdG}}=h(\mathbf{r})\mu^{3}+\Delta(\mathbf{r})\mu^{1}\,.$ (8)
Here $\mu^{i}$ is the Pauli matrices acting on the particle-hole space, and $h(\mathbf{r})$ is the qBM Hamiltonian in Eq. (1). The pairing amplitude $\Delta(\mathbf{r})$ is obtained self-consistently by decoupling a local attractive interaction and is spatially inhomogeneous due to quasiperiodicity [30].
Fig. 5 shows the average $\Delta$ and its distribution function for the qBM-BCS model at half-filling obtained with KPM-based mean-field theory [30]. Figs. 5(a) and (b) show the average order parameter $\Delta$ along the two cuts $A_{1}=0.4$ and $1$, respectively. Without the $A_{2}$ potential, the interaction strength is below the critical one for 2D Dirac fermions [42, 43, 44, 45] and $\Delta$ vanishes unless $A_{1}$ is very close to the magic angles. The superconductivity in the qBM model is not fine-tuned as in the BM model, and becomes nonzero as long as $A_{2}$ is not too weak. The strength of the superconductivity correlates with the change of the DOS at the Fermi energy, peaking near filaments with maximal DOS. However, the change in $\Delta$ is much shallower than that of the DOS. It results from the competition of the Fermi energy DOS and the spatial overlap of wave functions. In BCS theory, superconductivity benefits from a larger DOS at the Fermi energy. However, in the BCS-qBM model near the DOS peaks (the filaments), the spatial overlap of wave functions decays faster with energy difference than that in the lakes between filaments (not shown), which makes the wave-function enhancement of superconductivity less significant. Figs. 5(c) and (d) show the distribution of the local order parameter along the two cuts. We can see that the distribution peaks at nonzero values and vanishes at $\Delta_{i}=0$, indicating the superconductivity is nonzero almost everywhere.
[…]
Conclusion.—We have shown that the admixture of topology and quasiperiodicity in moiré materials can lead to remarkable quantum critical phenomena and robust superconductivity. Quasiperiodicity acts like disorder for strong mixing, but the coupling of same-valley Dirac fermions for small twists simulates topological surface states. The resulting wave function criticality can strongly boost superconducting pairing without magic-angle fine tuning and without degrading the superfluid stiffness.
While this result applies to twisted bilayer graphene in the chiral limit, we expect distinct multifractality could be found in trilayer systems as well as twisted bilayers with a superimposed, incommensurate superlattice. Even in symmetry classes where disorder localizes in two-dimensions, quasiperiodicity can drive transitions to critical states [48]; this opens up the possibility for similar physics to occur in other van der Waals heterostructures such as transition-metal dichalcogenides where the single-particle physics is simpler.
Acknowledgements.
We thank S. Gopalakrishnan for helpful discussions. This work was supported by the Welch Foundation Grant No. C-1809 (X. Z. and M. S. F.) and by the National Science Foundation under Grant No. DMR-2238895 (J. H. W.). This work was performed in part at Aspen Center for Physics, which is supported by NSF Grant No. PHY-2210452. This work was supported in part by the Big-Data Private-Cloud Research Cyberinfrastructure MRI-award funded by NSF under grant CNS-1338099 and by Rice University’s Center for Research Computing (CRC). Portions of this research were conducted with high performance computing resources provided by Louisiana State University (http://www.hpc.lsu.edu).