### v2

# Critical Filaments and Superconductivity in Quasiperiodic Twisted Bilayer Graphene

Abstract
Multilayer moiré materials can exhibit topological electronic features yet are inherently quasiperiodic—leading to wave function interference whose Anderson-localizing tendency can be mitigated by topology. We consider a quasiperiodic variant of the chiral Bistritzer-MacDonald model for twisted bilayer graphene with two incommensurate moiré potentials that serves as a toy model for twisted trilayer. We observe “filaments” linking magic angles with enhanced density of states and fractal wave functions that evade localization; states away from the filaments mimic fractal surface states of dirty topological superconductors. We demonstrate that topological quasiperiodicity can broadly enhance superconductivity without magic-angle fine-tuning.
Amplifying electronic interactions and uncovering topological phases drives much of the research in moiré materials [1]. Correlated phases including superconductivity were observed in twisted bilayer graphene (TBLG) [1], attributed to enhanced interactions due to flattening of the moiré bands near magic angles [2, 3, 4, 5]. Concurrently, topology is enabled in TBLG by the coupling of same-chirality Dirac electrons between the layers [7, 8, 6, 9].
Generic twist angles in moiré materials lead to incommensurate (quasiperiodic) tunneling, but this has negligible effects in TBLG. By contrast, quasiperiodicity is ubiquitous when multiple moiré patterns compete, as occurs in systems such as twisted trilayer graphene (TTLG), where superconductivity was recently observed [10, 11]. Previous work on quasiperiodic Dirac and moiré materials focused on tight-binding models with lattice incommensuration [12, 13, 14], and provided glimpses of a kind of wave-function criticality linked with “magic angles.” Wave-function criticality (multifractality) is typically associated to disorder-driven Anderson localization transitions [15]. Quasiperiodicity induces interference that can also drive Anderson localization, although the most-studied cases in one dimension (1D) show features different from random 1D systems [16, 17]. At the same time the tendency towards localization can be mitigated by topology, which protects boundary and bulk states in nontrivial bands.
The interplay between multifractality and interactions can produce surprising effects. For example, the average Cooper pairing amplitude can be enhanced near a superconductor-insulator transition if Coulomb interactions are externally screened. The mechanism involves the Russian-doll-like nesting (“Chalker scaling”) of multifractal states, and can occur near the localization transition [18, 19] and in 1D quasiperiodic models [20, 21, 22].
Figure 1: (a): Pairing amplitude $\Delta$ and density of states (DOS) $\nu(E)$ for the chiral Bistritzer-MacDonald (BM) model; the enhancement of superconductivity occurs near the magic-angle ($A_{1}\approx 0.586$). (b): The same for the quasiperiodic BM model; superconductivity is broadly enhanced by wave-function fractality without magic-angle fine-tuning. The parameters $A_{1,2}$ are the strengths of the two incommensurate BM potentials, see Eqs. (1) and (2). All results are obtained from self-consistent BCS numerics using the kernel polynomial method [23]; $U$ is the interaction strength measured relative to the kinetic energy. The inset shows the $U$-dependence of $\Delta$ at $A_{1}=1$ and $A_{2}=0.6$. Here the combination of Chalker scaling and topology [24, 25, 26] leads to BCS-like $\Delta\sim e^{-c/U}$ (see text), instead of the threshold behavior of a 2D Dirac semimetal. The latter occurs away from the magic angle in (a).
This raises a question: Can quasiperiodic and topological effects in multilayer moiré materials lead to criticality that enhances correlated behavior, such as superconductivity? We address this by studying a quasiperiodic variant of the chiral Bistritizer-MacDonald (BM) model for TBLG. Our model incorporates two incommensurate BM potentials as occurs in TTLG, and serves as a toy version of the latter with two layers instead of three. As we show in this Letter, quasiperiodicity with topology can lead to criticality and amplify pairing without the need to fine-tune to magic angles, see Fig. 1.
The BM model for TBLG [2, 4] possesses an emergent particle-hole symmetry [6], falling into class C in the $10$-fold classification of topological matter [27] and 2D Dirac Hamiltonians [28]. If AA tunneling vanishes, a chiral symmetry emerges [5]. This chiral BM model takes the same form as the 2D surface theory of a bulk class CI topological superconductor (TSC) [29], while the BM model corresponds to the same surface, subject to orbital time-reversal symmetry breaking.
Quenched disorder, which respects the chiral symmetry, leads to remarkable results for class-CI TSC surface states, including a power-law vanishing density of states (DOS) $\nu(E)\sim\left|E\right|^{1/7}$ [30] and critical low-energy wave functions with universal multifractal statistics [31]. Recent results have demonstrated that a more radical kind of universal wave-function criticality can extend over a wide range of the energy spectrum for these surfaces (“spectrum-wide quantum criticality,” SWQC [24, 25, 26]).
In this Letter, we show that topological quasiperiodicity leads to DOS enhancements and wave-function fractality. We establish that quasiperiodic multilayer graphene effectively simulates disordered TSC surface states. The resulting SWQC and absence of Anderson localization are exotic TSC features [24, 25] that are synthetically realizable in multilayer moiré materials. Our results are surprising because they precisely link the criticality of dirty surface states of bulk topological phases to quasiperiodic layered 2D materials. Since the bulk is “missing,” one might say that quasiperiodic twisted multilayers holographically realize bulk topological superconductor classes.
We study the consequences of topologically-protected quasiperiodicity for superconductivity using self-consistent BCS calculations [22], finding enhanced Cooper pairing [18, 19] without fine-tuning to magic angles (Fig. 1). For simplicity, we consider intravalley pairing. Our calculations incorporate additional aspects beyond those of the normal state such as Altshuler-Aronov corrections, which can also cause Anderson localization; such corrections shift the Bogoliubov-de Gennes eigenstate localization transition in a 1D quasiperiodic model [22]. We demonstrate that this effect is absent here due to topological protection [32], and we find that the superfluid stiffness precisely encodes the topological character of the normal state (another instance of the “holography”). Our approach can be applied to the more complicated case of TTLG [33].
Model.—We consider a quasiperiodic version of the chiral Bistritzer-MacDonald (qBM) model [4, 5],
$h(\mathbf{r})=-iv_{F}\,\bm{\sigma}\cdot\nabla+V_{\mathbf{q}}(\mathbf{r})+V_{\mathbf{q}^{\prime}}(\mathbf{r})\,.$ (1)
The Pauli matrices $\sigma_{1,2}$ act on the sublattice space, and $V_{\mathbf{q}}(\mathbf{r})$ and $V_{\mathbf{q}^{\prime}}(\mathbf{r})$ denote moiré potentials coupling the two layers in AB- and BA-stacking regimes,
[…]
Here $\tau^{\pm}=(\tau_{x}\pm i\tau_{y})/2$ (and similarly for $\sigma^{\pm}$) with $\tau_{j}$ as the Pauli matrices that link the layers. The potentials $U_{0}$ and $U_{1}$ take the forms [4] $U_{0}(\mathbf{r})=e^{-i\mathbf{q}_{1}\cdot\mathbf{r}}+e^{-i\mathbf{q}_{2}\cdot\mathbf{r}}+e^{-i\mathbf{q}_{3}\cdot\mathbf{r}}$ and $U_{1}(\mathbf{r})=e^{-i\mathbf{q}_{1}\cdot\mathbf{r}}+e^{-i\frac{2\pi}{3}}e^{-i\mathbf{q}_{2}\cdot\mathbf{r}}+e^{i\frac{2\pi}{3}}e^{-i\mathbf{q}_{3}\cdot\mathbf{r}}\,.$ Here $\mathbf{q}_{1}=k_{\theta}(1,0)$ and $\mathbf{q}_{2/3}=k_{\theta}(-1/2,\pm\sqrt{3}/2)$ are the reciprocal lattice vectors of the usual moiré potential. The momentum $k_{\theta}=2K\sin\theta/2$ defines the moiré scale (with $K$ the K-point wavevector modulus in graphene). The second moiré potential $V_{\mathbf{q}^{\prime}}(\mathbf{r})$ takes the same form but with different reciprocal vectors $\{\mathbf{q}^{\prime}_{i}\}$ and potential strength $A_{2}$. When the two sets $\{\mathbf{q}_{i}\}$ and $\{\mathbf{q}^{\prime}_{i}\}$ are mutually incommensurate, we have a quasiperiodic model. In this Letter, we focus on the case $\mathbf{q}_{i}=\beta\,\mathbf{q}_{i}^{\prime}$ with $\beta$ the inverse golden ratio $(\sqrt{5}-1)/2$; we have verified that similar results obtain for other (e.g., silver) ratios [23].
Figure 2: The zero-energy density of states $\nu(0)$ for the qBM model in chiral limit. Here $A_{1}$ and $A_{2}$ are the strengths of the two sets of moiré potentials. (a): The map of $\nu(0)$ for the two moiré potentials. The zero-energy DOS has large peaks at the magic angles of the $A_{1}$ and $A_{2}$ potentials [indicated by red stars in (a)]. The magic angles of $A_{1}$ and $A_{2}$ are connected by several filaments, where $\nu(0)$ peaks. (b), (c) and (d): $\nu(0)$ along the cuts at $A_{1}=0.4$, $1$ and $1.5$, respectively. The peaks correspond to the filaments and $\nu(0)$ remains nonzero between these. This is our first hint that quasi-periodicity has driven a single-particle phase transition away from the semimetal in these regions. The data are obtained with KPM for a momentum lattice with $N_{\Lambda}=17$ and $F_{n}=89$. We use Chebyshev expansion order $N_{\scriptscriptstyle\mathcal{C}}=2^{11}$ for the DOS map (a) and $N_{\scriptscriptstyle\mathcal{C}}=2^{11}$, $2^{12}$ and $2^{13}$ in (b-d).
The qBM model is realized in momentum space using a momentum cutoff $\Lambda=N_{\Lambda}k_{\theta}/2$. Focusing on $V_{\mathbf{q}}(\mathbf{r})$, the moiré Brillouin zone (mBZ) has linear size $k_{\theta}$, and we keep $2N_{\Lambda}^{2}$ moiré bands. Without a second potential, a momentum space lattice of linear size $N_{\Lambda}$ is sufficient; however, $\mathbf{q}^{\prime}_{i}$ allows hoppings within the mBZ. Since $|\mathbf{q}_{i}^{\prime}|=k_{\theta}/\beta$, we controllably reach larger system sizes with rational approximations of $\beta$, i.e. $\beta_{n}=F_{n}/F_{n+1}$ for Fibonacci numbers $F_{n}$. In this case, the denominator $\frac{|\mathbf{q}_{i}^{\prime}|}{k_{\theta}}=\frac{F_{n+1}}{F_{n}}$ determines the linear size of momentum space points required within one mBZ. Therefore, the full momentum space lattice has linear size $N_{\Lambda}F_{n}$ and spacing $\Delta k=k_{\theta}/F_{n}$; in the incommensurate limit ($n\rightarrow\infty$), all momentum can be accessed. In the following, we work in units such that $v_{F}=1$ and $k_{\theta}=\left|\mathbf{q}_{i}\right|=1$; all energies are measured in units of $v_{F}\,k_{\theta}$. Our setup is further illustrated in Ref. [23].
Density of states and fractal wave functions.—Dirac cones have a vanishing DOS $\nu(E)\sim|E|/v^{2}$ with velocity $v$; at the magic angle $v\rightarrow 0$ and $\nu(0)$ diverges. In the quasiperiodic case, at small $A_{1,2}$ the semimetal is stable and eigenstates remain plane-wave-like [34]. However, when $\nu(0)>0$ (or vanishes with a small power-law), multifractality–typically associated to an Anderson metal-insulator transition [15]–can occur. To reveal these (potential) single-particle phases, Fig. 2 shows $\nu(0)$ for the qBM model evaluated with the kernel polynomial method (KPM) [35]. Fig. 2(a) shows the $\nu(0)$ landscape while $A_{1,2}$ vary. There are large lakes with $\nu(0)>0$ separated by filaments where the DOS is strongly enhanced. The filaments connect the magic angles (red stars) of the monochromatic potentials; both filaments and lakes host multifractal wave functions (see below). Fig. 2(b-d) show particular cuts. In the regime close to the $A_{1}$ or $A_{2}$ axes away from the magic angles, the system remains semimetallic and $\nu(0)\rightarrow 0$. By contrast, $\nu(0)$ peaks at the filaments and either remains nonzero or vanishes with a nontrivial power-law between the filaments [23]. In the regions of comparable $A_{1}$ and $A_{2}$, between the filaments, $\nu(0)$ is generally nonzero; in fact, as we show below, the system exhibits SWQC behavior in these regions associated to class-CI surface states [24].
Figure 3: The second multifractal dimension $\tau_{2}$ averaged for states near zero energy and zero energy DOS $\nu(E=0)$ for the qBM model. (a) and (b): along the vertical cuts at $A_{1}=0.4$ and $1$, respectively; (c) and (d): along the horizontal cuts at $A_{2}=0.2$ and $A_{2}=0.5$, respectively. The navy blue dashed and orange dot-dashed lines indicate $\tau_{2}=3/2$ for the WZW states and $\tau_{2}=7/4$ for the SQHT states, respectively.
Both cutoff $\Lambda$ and number of moiré unit cells $F_{n}$ determine the momentum space lattice [23], which need to both be large; therefore, to determine wave functions, we employ the Chebyshev filtering and shaking method [36, 37]. This method obtains eigenfunctions in a small energy window, which could contain thousands of states near a narrow band [23]. We use the method to obtain eigenfunctions for the qBM model with $\sim 10^{6}$ lattice sites.
Figure 4: The second multifractal dimension $\tau_{2}$ versus energy and the wave functions at the one of filaments and in the lakes between the filaments. (a): $\tau_{2}$-$E$ at the filament with $A_{1}=0.4$ and $A_{2}=0.36$; (b): same as (a) but in the lakes between filaments with $A_{1}=0.4$ and $A_{2}=0.6$. (c)-(e): the wave functions at the filament for states at $E=0$, $0.2$ and $0.425$; (f)-(h): the wave functions in the lake for states at $E=0$, $0.2$ and $0.5$. The shadows in (a) and (b) show the energy gaps in the spectrum. The navy dashed and orange dot-dashed line in (a) and (b) show the reference value of $\tau_{2}$ for the WZW and SQHT states, respectively. Here we $N_{\Lambda}=11$, $F_{n}=55$ and the wave functions reside on a $605\times 605$ lattice in real space.
The statistics of single-particle wave functions can be characterized by the inverse participation ratio (IPR) $P_{q}\equiv\sum_{i}\left|\psi_{i}\right|^{2q}\,.$ Here $\psi_{i}$ is the (normalized) wave function in real space, obtained by Fourier transform of the momentum-space wave function. For extended wave functions, $P_{q}\sim L^{-d(q-1)}$, with $L$ the linear dimension of the system; for localized wave functions, $P_{q}\sim L^{0}$. Further, if $P_{q}\sim L^{-\tau_{q}}$ and $0<\tau_{q}<d(q-1)$, that wave function is critical. In the thermodynamic limit ($L\to\infty$), the multifractal dimension can be evaluated via $\tau_{q}=-\log{P_{q}}/\log{L}\,.$ The $\tau_{q}$ can also be obtained with the box-averaged IPR, $P_{q}^{\scriptscriptstyle\mathsf{box}}=\sum_{i_{b}}\left(\sum_{\text{box}}\left|\psi_{i}\right|^{2}\right)^{q}\propto b^{\tau_{q}}\,,$ with $b$ the linear size of the box. We use both methods to calculate the $\tau_{q}$ and they are consistent for the appropriate box sizes.
The second multifractal dimension $\tau_{2}$ for low-energy wave functions is shown in Fig. 3 for the qBM model along the cuts $A_{1}=0.4$ and $1$. Corresponding $\tau_{3}$ plots appear in [23]. The value $\tau_{2}\approx 2$ indicates ballistic wavefunctions, and Fig. 3(a) shows such states for small $A_{2}$ and $A_{1}=0.4$. We observe a correlation between $\nu(0)$ and multifractality; at the first filament (with increasing $A_{2}$), $\tau_{2}$ decreases, fluctuates for stronger $A_{2}$, and dips near the additional filaments. The value of $\tau_{2}$ near the filaments in Fig. 3(a,c) indicates that the wave functions there are strongly multifractal, see also Fig. 4(c,d,g). Note that Fig. 3(c) follows the pattern for magic-angle semimetals as outlined in [13, 14]. In other regimes, such as between the filaments, the wave functions are also multifractal with a different value of $\tau_{2}$. In the nonperturbative regime, e.g., Fig. 3(b) along $A_{1}=1$, $\tau_{2}$ fluctuates around $3/2$ and is only partially correlated with $\nu(0)$. However, at this point $\tau_{2}$ appears to have saturated to the known value $\tau_{2}=3/2$ for zero-energy Wess-Zumino-Witten (WZW) surface states of a disordered class-CI TSC [31, 15, 28]. [Fig. 3(d) also shows this saturation.]
Finite-energy surface-state wave functions of a dirty TSC can also evade localization via SWQC, with critical statistics tied to a quantum-Hall plateau transition [24, 38, 25, 26]. For the class-CI qBM model, we anticipate finite-energy states with SWQC associated to the spin quantum-Hall transition (SQHT) [24, 25]. To compare with finite energies in the qBM model, we take two separate values, on the filament ($A_{1}=0.4$ and $A_{2}=0.36$) and between the filaments ($A_{1}=0.4$ and $A_{2}=0.6$) [see Fig. 4(a,b)], and examine a range of energies, extracting $\tau_{2}$. On the filament [Fig. 4(a)], the low-energy states are strongly multifractal (with $\tau_{2}\approx 1$) with a few isolated peaks [energies $E=0,0.2$, Figs. 4(c,d)]; the higher energies oscillate between SQHT states [$\tau_{2}\approx 7/4$ and less strongly peaked, see Fig. 4(e)] and strongly multifractal states. The filaments, originating from the flat bands at the magic-angles, have hybridized those states strongly into these multifractal states. In between the filaments [Fig. 4(b)], the low-energy wave functions are less multifractal than on the filament with WZW-like states [$\tau_{2}\approx 3/2$ and $\tau_{3}\approx 5/2$, Fig. 4(g)] and the high-energy states that reside deep within moiré bands are SQHT states [$\tau_{2}\approx 7/4$ and $\tau_{3}\approx 13/4$, Fig. 4(h)]. Near the energy gaps, the wave functions show stronger fractal features.
[…]
$h_{\mathsf{BdG}}=h(\mathbf{r})\mu^{3}+\Delta(\mathbf{r})\mu^{1}\,.$ (3)
The Pauli matrices $\mu^{i}$ act on particle-hole space, and $h(\mathbf{r})$ is the qBM Hamiltonian in Eq. (1). The pairing amplitude $\Delta(\mathbf{r})$ is inhomogeneous due to quasiperiodicity and obtained self-consistently by decoupling a local attractive interaction with coupling strength $U$ [23].
Figure 5: The average superconductivity order parameter $\Delta$, normal-state zero-energy DOS $\nu(0)$, distribution of the local order parameter $\Delta_{i}$ and superfluid stiffness. (a): $\Delta$ and $\nu(E=0)$ along the cut $A_{1}=0.4$; (b): same as (a) along the cut $A_{1}=1$; (c): the distribution of local order parameter $\Delta_{i}$ along $A_{1}=0.4$; (d) same as (c) along the cut $A_{1}=1$; (e): superfluid stiffness (red) along $A_{1}=0.4$ (clean limit value $2\Delta/\pi$ indicated by black dashed curve); (f): same as (e) along $A_{1}=1$. Here $U=4v_{F}k_{\theta}$.
Fig. 5 shows the average $\Delta$ and its distribution function at half-filling obtained with KPM-based mean-field theory [23]. Figs. 5(a) and (b) plot the average order parameter $\Delta$ along the two cuts $A_{1}=0.4$ and $1$, respectively. Without the $A_{2}$ potential, the interaction strength is below the critical one for 2D Dirac fermions [39, 40, 41, 42] and $\Delta$ vanishes unless $A_{1}$ is close to the magic angles. Unlike the BM model, superconductivity in the qBM model is not fine-tuned and becomes nonzero as long as $A_{2}$ is not too weak. The strength of the superconductivity correlates with the change of the DOS at the Fermi energy, peaking near filaments with maximal DOS. However, the change in $\Delta$ is much shallower than that of the DOS. It results from the competition of the Fermi energy DOS and the spatial overlap of normal-state wave functions. In BCS theory, superconductivity benefits from a larger DOS at the Fermi energy. Here however, near the DOS peaks (the filaments), the spatial overlap of wave functions decays faster with energy difference than that in the lakes between filaments (not shown), which reduces the enhancement of superconductivity. Figs. 5(c) and (d) show the distribution of $\Delta_{i}$ and indicate that superconductivity is nonzero almost everywhere.
We compare $\Delta$ in the qBM and usual chiral BM models in Fig. 1. Fig. 1(b) shows $\Delta$ as a function of $U$ in the qBM model for a particular $A_{1,2}$ choice. The scaling indicates BCS-like behavior $\Delta\sim\exp(-c/U)$, where $c$ is a constant, in the $U\rightarrow 0$ limit. This should be contrasted with the chiral BM model, where $\Delta=0$ identically for sufficiently small $U$ away from the magic angle. In [23], we show that finite-energy SQHT states (present due to SWQC) and critical DOS scaling for class CI implies BCS scaling.
[…]
Conclusion.—We have shown that the admixture of topology and quasiperiodicity in moiré materials can lead to remarkable quantum critical phenomena and robust superconductivity. Quasiperiodicity acts like disorder for strong mixing, but the coupling of same-valley Dirac fermions for small twists simulates topological surface states. Superconductivity is boosted without magic-angle fine tuning and without degrading the superfluid stiffness.
Taking this toy model seriously, it is worth pointing out that in the microscopic TBLG lattice, our pairing is formally Fulde–Ferrell–Larkin–Ovchinnikov, which would imply a pair-density wave on the lattice scale; further, it would be realized along with the pairing in the other valley to preserve time-reversal. However, the chiral, quasiperiodic BM model we use in Eq. (1) can be exactly constructed in a single layer of graphene with two mismatched $\sqrt{3}$-potentials, as was shown for one such potential in Ref. [45]; in this system, pairing occurs between valleys (and is therefore, s-wave).
The qBM model is a toy version of TTLG, where we expect distinct multifractality will be found [33], as well as in twisted bilayers with a superimposed, incommensurate superlattice. Topology plays an important role; while quasiperiodicity can drive transitions to critical states [46], it may not enhance and could even destroy superconductivity [47]. Nonetheless, this opens up the possibility for similar physics to occur in other moiré materials such as transition-metal dichalcogenide multilayers.
Acknowledgements.
We thank S. Gopalakrishnan and E. König for helpful discussions. We further thank Aaron Dunbrack for pointing out the physical system that realizes our model. This work was supported by the Welch Foundation Grant No. C-1809 (X. Z. and M. S. F.) and by the National Science Foundation under Grant No. DMR-2238895 (J. H. W.). This work was performed in part at Aspen Center for Physics, which is supported by NSF Grant No. PHY-2210452. This work was supported in part by the Big-Data Private-Cloud Research Cyberinfrastructure MRI-award funded by NSF under grant CNS-1338099 and by Rice University’s Center for Research Computing (CRC). Portions of this research were conducted with high performance computing resources provided by Louisiana State University (http://www.hpc.lsu.edu).